OpenStudy (anonymous):

For a recent job, an electrician earned $50 per hour, and the electrician's apprentice earned $20 per hour. The electrician worked 6 hours more than the apprentice, and together they earned a total of $650. How much did each person earn? I know the electrician earned $550 and his apprentice earned $100 by other using another method, but i have to make the 2 equations that help solve this problem.

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

50e + 20a = 650 e = a + 6 I think this is right.

OpenStudy (anonymous):

e = a + 6 50(a+6) + 20a = 650 50a + 300 + 20a = 650 70a = 350 a = 5(hours) 5 * 20 = 100 5+6 = 11 11*50 = 550 Maybe not exactly.

OpenStudy (anonymous):

well in know the electrician earned 550 and the apprentice earned 100 therefore the electrician worked 11 hrs and the apprentice worked 5 hrs

OpenStudy (anonymous):

are those the equations you need?

OpenStudy (anonymous):

Yes, i think these are it. Thank you.